OpenStudy (anonymous):

A certain store has a fax machine available for use by its customers. The store charges $2.30 to send the first page and $0.40 for each subsequent page. The total price, P, for the faxing x pages can be modeled by the formula . Determine the number of pages that can be faxed for $5.90.

OpenStudy (anonymous):

P (price) = 2.30 + (0.40) * x x represents the subsequent pages

OpenStudy (anonymous):

so, 5.90 = 2.30 + (0.40) * x solve for x

OpenStudy (anonymous):

5.90 - 2.30 = 0.40 * x

OpenStudy (anonymous):

4.60 / 0.40 = x

OpenStudy (anonymous):

sorry thats 3.60 not 4.60 / 0.40

OpenStudy (anonymous):

so x = 9 subsequent pages but add 1 for the first page that is 2.30 so total of 10 pages